A chemistry student need 15L of a 32% acid solution. The laboratory has ample amounts of 30% and 40% acid solutions. How many liters of each solution should the student mix?
Answer by mananth(16946)

percent ---------------- quantity
Acid solution I 30 ---------------- x L
Acid solution II 42 ------ 15 - x L
Mixture 32.00% ---------------- 15

30 x + 42 ( 15 - x ) = 32 * 15

30 x + 630 - 42 x = 480
30 x - 42 x = 480 - 630
-12 x = -150
/ -12
x= 12.5L 30.00% Acid solution I
2.5L 42.00% Acid solution II

m.ananth@hotmail.ca